OpenStudy (anonymous):

roberto needs 6 cookies and 2 brownes for every 4 plates how many cookies and brownies will he need for 10 plates

OpenStudy (anonymous):

18 c 6 b

OpenStudy (anonymous):

there has to be enough for all 10 plates so you have to go over. so 18 cookies and 6 brownies.

OpenStudy (anonymous):

okay so you want to find out how many brownies is needed for 1 plate and how many cookies is needed for one plate. So to find out how many brownies is needed for one plate you devide the number of brownies by the amount of plates it satisfies and do the same for cookies. so 6/4=1.5 so 1.5 brownies per plate as well as 2/4=.5 so half a cookie per plate. now multiply each by the number of plates you are curious for and you will get your answer: 1.5*10=15 so 15 brownies per plate and .5*10=5 cookies per plate

OpenStudy (anonymous):

I mixed up my cookies and brownies, it would be 15 cookies and 5 brownies!